An electron is emitted from the surface of the big metallic plane perpendicular to it with initial velocity 0 v , How far from the plane the electron may manage to move in two cases: 1) the electron moves in the braking electric field E , 2) additionally to the electric field E, there is a magnetic field B, directed parallel to the surface of the plane
1
Expert's answer
2020-01-20T05:24:17-0500

1)

the kinetic energy of the electron is numerically equal to the work of the forces of the electric field

"\frac{mv^2}{2}=eEd (1)"

Where m is the mass of electron, e is the charge of electron

Using (1) we get

"d=\frac{mv^2}{eE}"


2) the magnetic force is always perpendicular to the direction of displacement and magnetic forces do not do any work

"d=\frac{mv^2}{eE}"